Given that P(A)=0.8, P(A/B)=0.8 , P(A∩B)=0.5, Find


(i)P(A/AUB)  Answer-32/37


(ii)P(A∩B/AUB)  Answer-20/37

Grade:12

2 Answers

jitender lakhanpal
12 years ago

Dear Shaleen,

  P(A)=0.8, P(A/B)=0.8 , P(A∩B)=0.5       Given

and we know that the conditional probability 

P(A/B)= P(A∩B)/ P(B)   so we can find  P(B)= 5/8

from this we can find  P(AUB) = P(A)+P(B)-P(A∩B)

P(AUB) = 37/40

P(A/AUB) =  P(A(AUB)/ P(AUB)

as     P(A(AUB) =P(A)    set theory 

so we get 

 P(A/AUB) = P(A)/ P(AUB)

 P(A/AUB)=(8/10)/(37/40)

 P(A/AUB)=32/37  

2)       

P(A∩B/AUB)= ( P(A∩B)(AUB))/ P(AUB)

P(A∩B)(AUB) =P(A∩B)

P(A∩B/AUB)=  P(A∩B)/ P(AUB)

putting the values we get 

P(A∩B/AUB)= 20/37  



We are all IITians and here to help you in your IIT JEE preparation.

Now you can win by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar : Click here to download the toolbar..  

jitender

askiitians expert

Ashwin Muralidharan IIT Madras
12 years ago

Hi Shaleen,

 

P(A)=0.8, P(A/B)=0.8 , P(A∩B)=0.5       

 

Now conditional Probability,

P(A/B)= P(A∩B)/ P(B)   

Hence  P(B)= 5/8

 

Next, P(AUB) = P(A)+P(B)-P(A∩B)

So P(AUB) = 37/40

 

And, P(A/AUB) =  P(A∩(AUB)/ P(AUB) -------[as P(A∩(AUB) =P(A)]

So,

 P(A/AUB) = P(A)/ P(AUB)

 P(A/AUB)=(8/10)/(37/40)

 P(A/AUB)=32/37  

 

 For the Next one,

P(A∩B/AUB)= ( P(A∩B)∩(AUB))/ P(AUB)

P(A∩B)∩(AUB) =P(A∩B)

 P(A∩B/AUB)=  P(A∩B)/ P(AUB)

Substituting the values from above,

P(A∩B/AUB)= 20/37  

That solves the two parts.

 

Hope that helps.

 

Best Regards,

Ashwin (IIT Madras).